# February 12, 2024

### v1

@@ II.1 Scalar Sector
For the potential of the $Z_{3}$ 3HDM model, denoted $V_{Z_{3}}$, we use the conventions of [6]. The $Z_{3}$-invariant terms, given by $\phi_{i}\to\phi_{i}^{\prime}=(S_{Z_{3}})_{ij}\phi_{j}$, can be expressed as:
@@
$\begin{pmatrix}H_{0}\\ R_{1}\\ R_{2}\end{pmatrix}=\mathcal{O}_{\beta}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}c_{\beta_{2}}c_{\beta_{1}}&c_{\beta_{2}}s_{\beta_{1}}&s_{\beta_{2}}\\ -s_{\beta_{1}}&c_{\beta_{1}}&0\\ -c_{\beta_{1}}s_{\beta_{2}}&-s_{\beta_{1}}s_{\beta_{2}}&c_{\beta_{2}}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix},$ (6)
where here we have used the short notation, $c_{x}\equiv\cos x,s_{x}\equiv\sin x$.
@@
$\mathcal{R}_{1}=\begin{pmatrix}c_{\alpha_{1}}&s_{\alpha_{1}}&0\\ -s_{\alpha_{1}}&c_{\alpha_{1}}&0\\ 0&0&1\end{pmatrix}\,,\quad\mathcal{R}_{2}=\begin{pmatrix}c_{\alpha_{2}}&0&s_{\alpha_{2}}\\ 0&1&0\\ -s_{\alpha_{2}}&0&c_{\alpha_{2}}\end{pmatrix}\,,\quad\mathcal{R}_{3}=\begin{pmatrix}1&0&0\\ 0&c_{\alpha_{3}}&s_{\alpha_{3}}\\ 0&-s_{\alpha_{3}}&c_{\alpha_{3}}\end{pmatrix}\,.\quad$ (9)
@@
$\mathcal{O}_{\gamma_{1}}=\begin{pmatrix}1&0&0\\ 0&c_{\gamma_{1}}&-s_{\gamma_{1}}\\ 0&s_{\gamma_{1}}&c_{\gamma_{1}}\end{pmatrix},\qquad\mathcal{O}_{\gamma_{2}}=\begin{pmatrix}1&0&0\\ 0&c_{\gamma_{2}}&-s_{\gamma_{2}}\\ 0&s_{\gamma_{2}}&c_{\gamma_{2}}\end{pmatrix}.$ (11)
@@ III Constraints
$m_{C1}=m_{A1}\,,\qquad m_{C2}=m_{A2}\,,\qquad\gamma_{1}=\gamma_{2}\,.$ (19)
@@ IV Random Scan strategy
$\alpha_{1}\,,\alpha_{2}\,,\alpha_{3}\,,\gamma_{1}\,,\gamma_{2},\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right];\qquad\tan{\beta_{1}}\,,\tan{\beta_{2}},\in\left[0.3,10\right];$
$m_{H_{1}}\equiv m_{h_{2}}\,,m_{H_{2}}\equiv m_{h_{3}},\in\left[125,1000\right],\text{GeV};$ (20)
$m_{A_{1}}\,,m_{A_{2}},m_{H_{1}^{\pm}}\,,m_{H_{2}^{\pm}},\in\left[100,1000\right],\text{GeV};$ (21)
$m^{2}_{12},m^{2}_{13},m^{2}_{23}\in\left[\pm 10^{-1},\pm 10^{7}\right],\text{GeV}^{2}\,,$ (22)
where the last expression applies only to the soft masses that are not obtained as derived quantities (see Ref.[10] for the complete expressions).
@@
$\gamma_{1}=\gamma_{2}=-\alpha_{3},\quad m_{H_{1}}=m_{A_{1}}=m_{H_{1}^{\pm}},\quad m_{H_{2}}=m_{A_{2}}=m_{H_{2}^{\pm}},$ (23)
@@ V.3 Further Implementation Details
Furthermore, we follow the methodology in [27] where we define all operations related to CMAES and density estimation not over the parameter space, but over a hypercube $[0,1]^{d_{\mathcal{P}}}$, where $d_{\mathcal{P}}$ is the number of parameters, that maps to the parameter space $\mathcal{P}$. This allows us to treat all parameters in an equal nominal footing, avoiding any potential pathologies arising from having different parameters spanning many different orders of magnitude.
@@ VI Analysis and results
We now present and analyse the results for multiple scans using the ideas presented in section V, and two random sampling scan strategies: purely random over the whole parameter space and $50\%$ away from the alignment limit defined by eq. 27. We present two analysis, one where the HiggsBounds constraints using HiggsTools were not included in the loss function, and one where it has. The scan without HiggsTools runs faster in both computational overhead and CMAES convergence (to be discussed below), which allowed us to experiment with our methodology. The impact of HiggsTools on points obtained without including it in the loop is studied. We then perform a second analysis where we included HiggsTools in the loop to show how one can include multiple constraints from different sources and still be able to explore the whole parameter space of the model.
@@
Sampling Scan $d_{\mathcal{P}}$ $N_{C}$ $p_{\mathcal{P}}$ $p_{\mathcal{O}}$ Random Completely random $16$ $61$ N/A N/A Around alignment: AL-2 $16$ $61$ N/A N/A CMAES No penalty (Vanilla) $16$ $61$ None None Parameter space penalty $16$ $61$ $16$ parameters None Constraint space penalty $16$ $61$ None $61$ quantities $\mathcal{O}_{i}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus $16$ $61$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ None $\mu_{ggF,\gamma\gamma}$ and $\mu_{ggF,Z\gamma}$ focus $16$ $61$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ $m_{H_{1}^{+}}$ and $m_{H_{2}^{+}}$ $16$ $61$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ None CMAES Parameter space penalty $16$ $67$ $17$ parameters None w/ HT Constraint space penalty $16$ $67$ None $67$ quantities $\mathcal{O}_{i}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus $16$ $67$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ None $\mu_{ggF,\gamma\gamma}$ and $\mu_{ggF,Z\gamma}$ focus $16$ $67$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ $m_{H_{1}^{+}}$ and $m_{H_{2}^{+}}$ $16$ $67$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ None
@@ VI.1 Rewarding Exploration in the Parameter Space
We first study the implementation of CMAES with and without parameter space reward to show the enhanced exploration capabilities of CMAES when including a parameter density penalty in the loss function. In fig. 1 we show the scatter plot of the $(\sin(\alpha_{1}-\beta_{1}),\sin(\alpha_{2}-\beta_{2}))$ plane for different runs. In particular, we exhibit the difficulty of random sampling finding valid points, with fig. 0(a) showing only 23 valid points, of which only one passed the HiggsBounds constraints. These were obtained from a scan that sampled an estimated $\mathcal{O}(10^{13})$. [^7] We can only estimate the number of points as it would have been prohibitive to store all non-valid points. Therefore, we measured how long the scan took to process a few thousand points, and kept a loose track of the random sampling run to produce the estimate. In fig. 0(b) we show the points obtained by sampling within 50% of the alignment limit, where the allowed points are highly constrained with $\alpha_{1}\simeq\beta_{1}$, an expected result due to the highly constraining bounds from collider measurements of the Standard Model-like Higgs boson decay channels. In the same graph, we can also observe the boundaries imposed by the alignment limit, as $|\sin(\alpha_{2}-\beta_{2})|\simeq 0.5$. In the next plot, fig. 0(c), we show the result of the CMAES scan without further exploration. We observe a funnelling of the results into a single region $\alpha_{1},\ \beta_{1},\ \alpha_{2},\ \beta_{2}\simeq 0$, clearly providing little coverage of the parameter space, although still providing far more valid points than the random sampler. This lack of exploration of CMAES was first observed in [27] and is easily understood from an algorithm point of view as CMAES does not have built-in mechanisms to escape a minimum (global or minimal). [^8] In [27] CMAES was endowed with a restart strategy, which mitigates this and allowed CMAES to draw a more complete picture of the valid regions of the parameter space. In this paper, we have not implemented this, as our focus is on developing a novelty reward driven exploration.
@@
(b) Alignment limit sampling
(c) CMAES without exploration
@@
(e) CMAES with parameter novelty reward focused on $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$
@@
(b) CMAES with parameter novelty reward focused on $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$
@@
(b) Alignment limit sampling
(c) CMAES without exploration
@@
(e) CMAES with parameter novelty reward focused on $(m_{H^{+}_{1}},m_{H^{+}_{2}})$
@@
In fig. 2(d) and fig. 2(e) we show the results of the scans with the parameter density penalty over all parameters and focused on the charged masses, respectively. We observe that both were able to cover more parameter space than the alignment limit random scan, but the scan focused on the charged scalar masses was able to cover more of the $(m_{H^{+}_{1}},m_{H^{+}_{2}})$ plane, especially in the $m_{H^{+}_{1/2}}\gtrsim 100$ GeV limits. This result further shows that focussing the exploration reward on subsets of parameters can help uncover novel regions overlooked by traditional scans, although in this case most of the points with $m_{H^{+}_{1/2}}\gtrsim 100$ GeV did not survive HiggsBounds.
@@
(b) CMAES with parameter novelty reward focused on $(m_{H^{+}_{1}},m_{H^{+}_{2}})$
@@ VI.2 Rewarding Exploration in the Observable Space
(b) Alignment limit sampling
(c) CMAES without exploration
@@
(b) CMAES with observable novelty reward focused on $(\mu_{ggF,\gamma\gamma},\mu_{ggF,ZZ})$
@@ VI.4 Convergence Metrics
Figure 12: Total loss as function of generation. Only the first 500 generations are shown. The random sampler curves are over random generations of $12$ poitns, the same population size as CMAES. The shaded regions represent $0.95$ confidence intervals computed using a bootstrap of 100 runs.
@@
Despite the suggestive previous plot, not all scans converge within the budget, which we set to 100 runs of 2000 generations for each case in table 1 without HiggsTools, and to 200 runs of 2000 generations for the cases with HiggsTools. [^12] One could alternatively increase the budget for the HiggsTools by increasing the number of generations to 4000. Intuitively, more runs provide a more global picture, whereas longer runs allow for longer explorations of the valid region. The choice between allocating more budget to one over the other depends on the intended study. We present these metrics in table 2 alongside statistics on the fraction of valid points that are within the alignment cases, AL-1 from eq. 26 and AL-2 from eq. 27. We see that, while most points are within AL-1, only a minority are within AL-2. More interestingly, we observe how the scan with novelty reward in the $\alpha_{1},\ \beta_{1},\ \alpha_{2},\ \beta_{2}$ subspace of the parameter space has produced the most points away from the alignment limit. This can be visually understood in fig. 1 and fig. 2 where it is clear that the novelty reward is guiding CMAES to values of $\alpha_{1},\ \beta_{1}$ that are beyond the alignment limit bounds. This is a feature of the versatility of our methodology, as we can perform dedicated scans explicitly away from previously considered priors and regions of the parameter space.
@@
Sampling Scan Converged Runs Within AL-1 Within AL-2 CMAES No penalty (Vanilla) 97 out of 100 $1.00$ $1.5\times 10^{-2}$ Parameter space penalty 95 out of 100 $0.90$ $5.1\times 10^{-3}$ Constraint space penalty 90 out of 100 $0.99$ $1.8\times 10^{-4}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus 94 out of 100 $0.85$ $5.9\times 10^{-3}$ $\mu_{ggF,\gamma\gamma}$ and $\mu_{ggF,Z\gamma}$ focus 91 out of 100 $1.00$ $4.2\times 10^{-5}$ $m_{H_{1}^{+}}$ and $m_{H_{2}^{+}}$ 90 out of 100 $0.95$ $6.6\times 10^{-3}$ CMAES Parameter space penalty 92 out of 200 $0.92$ $1.1\times 10^{-2}$ w/ HT Constraint space penalty 111 out of 200 $0.94$ $1.4\times 10^{-2}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus 102 out of 200 $0.84$ $4.1\times 10^{-2}$ $\mu_{ggF,\gamma\gamma}$ and $\mu_{ggF,Z\gamma}$ focus 101 out of 200 $0.93$ $4.1\times 10^{-3}$ $m_{H_{1}^{+}}$ and $m_{H_{2}^{+}}$ 91 out of 200 $0.98$ $5.3\times 10^{-2}$
@@
Table 2: Convergence and coverage statistics of the different scans presented in table 1. While the first CMAES rows correspond to scans without HiggsTools in the loop, the corresponding fractions of points in both alignment cases are computed using points that have passed HiggsBounds constraints after the scan.
@@
Figure 13: Number of valid points founds as function of generation. Only the first 500 generations are shown. The random sampler curves are over random generations of $12$ points, the same population size as CMAES. The shaded regions represent $0.95$ confidence intervals computed using a bootstrap of 100 runs.
Before HT After HT Points Efficiency Points Efficiency AL-1 21 in $4.4\times 10^{12}$ $\mathcal{O}(10^{-11})$ 0 in $4.4\times 10^{12}$ $<\mathcal{O}(10^{-12})$ AL-2 13 701 in $10^{10}$ $\mathcal{O}(10^{-6})$ 510 in ×$\times 10^{10}$ $\mathcal{O}(10^{-8})$ Random 23 in $10^{13}$ $\mathcal{O}(10^{-12})$ 1 in $10^{13}$ $\mathcal{O}(10^{-13})$
@@
In fig. 14 we show the distribution of the generation where the first valid point was valid for the runs with and without HiggsTools in the loop. We see that checking for HiggsBounds postpones the discovery of the first good point by a factor of around $2$ in terms of generation number. However, this is far better than with random sampling (both purely random and around the alignment limit) where only at most $5\%$ of the points that respect all other constraints survive HiggsBounds constraints, table 3. Additionally, we note that the number of trial points needed to find a good valid point is around $\mathcal{O}(250-750)\times 12$ when not including HiggsBounds constraints in the loop, and $\mathcal{O}(400-1300)\times 12$ when including HiggsBounds constraints in the loop. Recall that the efficient sampling of the random sampler is (estimated) to be $\mathcal{O}(10^{-12})$ and $\mathcal{O}(10^{-13})$ respectively, which means that our methodology improves the sampling efficiency by a factor of 8 orders of magnitude, even if we do not consider the near perfect sampling efficiency after convergence during the density penalty-guided exploration phase. Unsurprisingly, the improvement over the AL-2 alignment limit sampling strategy is more modest, with around four orders of magnitude when considering HiggsBounds constraints, but we reiterate that this only pertains to the first convergence and that our algorithm can then continue to explore the region found and go beyond the alignment limit bounds, as was highlighted in the previous section.

### v2

# March 7, 2025

@@ II.1 Scalar Sector
For the potential of the $Z_{3}$ 3HDM model, denoted $V_{Z_{3}}$, we use the conventions of [6]. The $Z_{3}$-invariant terms, given by $\phi_{i}\to\phi_{i}^{\prime}=(S_{Z_{3}})_{ij}\phi_{j}$, where
$S_{Z_{3}}=\text{diag}(1,e^{i\frac{2\pi}{3}}e^{-i\frac{2\pi}{3}}).$ (1)
can be expressed as:
@@
$\begin{pmatrix}H_{0}\\ R_{1}\\ R_{2}\end{pmatrix}=\mathcal{O}_{\beta}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}=\begin{pmatrix}\cos{\beta_{2}}\cos{\beta_{1}}&\cos{\beta_{2}}\sin{\beta_{1}}&\sin{\beta_{2}}\\ -\sin{\beta_{1}}&\cos{\beta_{1}}&0\\ -\cos{\beta_{1}}\sin{\beta_{2}}&-\sin{\beta_{1}}\sin{\beta_{2}}&\cos{\beta_{2}}\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\end{pmatrix}.$ (7)
@@
$\mathcal{R}_{1}=\begin{pmatrix}\cos{\alpha_{1}}&\sin{\alpha_{1}}&0\\ -\sin{\alpha_{1}}&\cos{\alpha_{1}}&0\\ 0&0&1\end{pmatrix},\,\mathcal{R}_{2}=\begin{pmatrix}\cos{\alpha_{2}}&0&\sin{\alpha_{2}}\\ 0&1&0\\ -\sin{\alpha_{2}}&0&\cos{\alpha_{2}}\end{pmatrix},\,\mathcal{R}_{3}=\begin{pmatrix}1&0&0\\ 0&\cos{\alpha_{3}}&\sin{\alpha_{3}}\\ 0&-\sin{\alpha_{3}}&\cos{\alpha_{3}}\end{pmatrix}.$ (10)
@@
$\mathcal{O}_{\gamma_{1}}=\begin{pmatrix}1&0&0\\ 0&\cos{\gamma_{1}}&-\sin{\gamma_{1}}\\ 0&\sin{\gamma_{1}}&\cos{\gamma_{1}}\end{pmatrix},\qquad\mathcal{O}_{\gamma_{2}}=\begin{pmatrix}1&0&0\\ 0&\cos{\gamma_{2}}&-\sin{\gamma_{2}}\\ 0&\sin{\gamma_{2}}&\cos{\gamma_{2}}\end{pmatrix}.$ (12)
@@ III Constraints
$m_{H_{1}^{\pm}}=m_{A1}\,,\qquad m_{H_{2}^{\pm}}=m_{A2}\,,\qquad\gamma_{1}=\gamma_{2}\,.$ (20)
@@ IV Random Scan strategy
$\alpha_{1}\,,\alpha_{2}\,,\alpha_{3}\,,\gamma_{1}\,,\gamma_{2}\,\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right];\qquad\tan{\beta_{1}}\,,\tan{\beta_{2}}\,\in\left[0.3,10\right];$
$m_{H_{1}}\equiv m_{h_{2}}\,,m_{H_{2}}\equiv m_{h_{3}}\,\in\left[125,1000\right]\,\text{GeV};$ (21)
$m_{A_{1}}\,,m_{A_{2}},m_{H_{1}^{\pm}}\,,m_{H_{2}^{\pm}}\,\in\left[100,1000\right]\,\text{GeV};$ (22)
$m^{2}_{12},m^{2}_{13},m^{2}_{23}\in\left[\pm 10^{-1},\pm 10^{7}\right]\,\text{GeV}^{2}\,,$ (23)
where the last expression applies only to the soft masses that are not obtained as derived quantities (see Ref.[10] for the complete expressions). Note that $m^{2}_{ij}$ have dimensions of $(\text{Mass})^{2}$, but despite their notation they are not positive definite.
@@
$\gamma_{1}=\gamma_{2}=-\alpha_{3},\quad m_{H_{1}}=m_{A_{1}}=m_{H_{1}^{\pm}},\quad m_{H_{2}}=m_{A_{2}}=m_{H_{2}^{\pm}}.$ (24)
@@ V.3 Further Implementation Details
Furthermore, we follow the methodology in [28] where we define all operations related to CMAES and density estimation not over the parameter space, but over a hypercube $[0,1]^{d_{\mathcal{P}}}$, where $d_{\mathcal{P}}$ is the number of parameters, that maps to the parameter space $\mathcal{P}$. This allows us to treat all parameters in an equal nominal footing, avoiding any potential pathologies arising from having different parameters spanning many different orders of magnitude. For all scans the parameter space bounds are the same used for the random scans, defined in eq. 23.
@@ VI Analysis and results
We now present and analyse the results for multiple scans using the ideas presented in section V, and two random sampling scan strategies: purely random over the whole parameter space and $50\%$ away from the alignment limit defined by eq. 28. We present two analysis, one where the HiggsBounds constraints using HiggsTools were not included in the loss function, and one where it has. The scan without HiggsTools runs faster in both computational overhead and CMAES convergence (to be discussed below), which allowed us to experiment with our methodology. The impact of using HiggsTools, a posteriori, on points obtained without including it in the loop is studied. We then perform a second analysis where we included HiggsTools in the loop to show how one can include multiple constraints from different sources and still be able to explore the whole parameter space of the model.
@@
Sampling Scan $d_{\mathcal{P}}$ $N_{C}$ $p_{\mathcal{P}}$ $p_{\mathcal{O}}$ Random Completely random $16$ $61$ N/A N/A Alignment limit: AL-2 $16$ $61$ N/A N/A CMAES No penalty $16$ $61$ None None Parameter novelty reward $16$ $61$ $16$ parameters None Observable novelty reward $16$ $61$ None $61$ quantities $\mathcal{O}_{i}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus $16$ $61$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ focus $16$ $61$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ focus $16$ $61$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ None CMAES Parameter novelty reward $16$ $67$ $17$ parameters None w/ HiggsTools Observable novelty reward $16$ $67$ None $67$ quantities $\mathcal{O}_{i}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus $16$ $67$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ focus $16$ $67$ None $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ focus $16$ $67$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ None
@@ VI.1 Rewarding Exploration in the Parameter Space
We first study the implementation of CMAES with and without parameter space reward to show the enhanced exploration capabilities of CMAES when including a parameter density penalty in the loss function. In fig. 1 we show the scatter plot of the $(\sin(\alpha_{1}-\beta_{1}),\sin(\alpha_{2}-\beta_{2}))$ plane for different runs. In particular, we exhibit the difficulty of random sampling finding valid points, with fig. 1(a) showing only 23 valid points, of which only one passed the HiggsBounds constraints. These were obtained from a scan that sampled an estimated $\mathcal{O}(10^{13})$. [^7] We can only estimate the number of points as it would have been prohibitive to store all non-valid points. Therefore, we measured how long the scan took to process a few thousand points, and kept a loose track of the random sampling run to produce the estimate. In fig. 1(b) we show the points obtained by sampling within 50% of the alignment limit, where the allowed points are highly constrained with $\alpha_{1}\simeq\beta_{1}$, an expected result due to the highly constraining bounds from collider measurements of the Standard Model-like Higgs boson decay channels. In the same graph, we can also observe the boundaries imposed by the alignment limit, as $|\sin(\alpha_{1}-\beta_{1})|\simeq 0.5$ and how $\alpha_{2}$ are $\beta_{2}$ are considerably more constrained leading to $|\sin(\alpha_{2}-\beta_{2})|\simeq 0.25$. In the next plot, fig. 1(c), we show the result of the CMAES scan without further exploration. We observe a funnelling of the results into a single region $\alpha_{1},\ \beta_{1},\ \alpha_{2},\ \beta_{2}\simeq 0$, clearly providing little coverage of the parameter space, although still providing far more valid points than the random sampler. This lack of exploration of CMAES was first observed in [28] and is easily understood from an algorithm point of view as CMAES does not have built-in mechanisms to escape a minimum (global or minimal). [^8] In [28] CMAES was endowed with a restart strategy, which mitigates this and allowed CMAES to draw a more complete picture of the valid regions of the parameter space. In this paper, we have not implemented this, as our focus is on developing a novelty reward driven exploration.
@@
(b) Alignment limit: AL-2
(c) CMAES no penalty
@@
(e) CMAES with $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}$ focus
@@
(b) CMAES with $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}$ focus
@@
(b) Alignment limit: AL-2
(c) CMAES no penalty
@@
(e) CMAES with $m_{H^{+}_{1}},m_{H^{+}_{2}}$ focus
@@
In fig. 3(d) and fig. 3(e) we show the results of the scans with the parameter density penalty over all parameters and focused on the charged masses, respectively. We observe that both were able to cover more parameter space than the alignment limit random scan, but the scan focused on the charged scalar masses was able to cover more of the $(m_{H^{+}_{1}},m_{H^{+}_{2}})$ plane, especially in the $m_{H^{+}_{1/2}}\gtrsim 100$ GeV limits. This result further shows that focussing the exploration reward on subsets of parameters can help uncover novel regions overlooked by traditional scans, although in this case most of the points with $m_{H^{+}_{1/2}}\gtrsim 100$ GeV did not survive HiggsBounds. Furthermore, considering that all scans presented in fig. 3 share the same parameter boundaries, c.f. eq. 23, these results highlight how our methodology can explore regions uncovered by the low sampling efficiency of random sampler and by the simplified priors of the alignment limit.
In fig. 3 we can also observe a shortcoming of the method which we will encounter throughout this work. We can see that different runs lead to different sets of valid points, preventing us from drawing a global picture of the allowed points. This is due to the fact that, in each run CMAES traverses a very localised path in the parameter space until either the maximum number of generations has been reached or a local minimum stopping criterion is triggered. To mitigate this, we implemented each scan as a collection of multiple runs, but the problem has not been completely resolved. In sections VI.1.1 and VI.3 we will present refinements of the methodology that will allow us to better fill in the voids left by these scans.
@@
(b) CMAES with $m_{H^{+}_{1}},m_{H^{+}_{2}}$ focus
@@ VI.2 Rewarding Exploration in the Observable Space
(b) Alignment limit: AL-2
(c) CMAES no penalty
@@
(b) CMAES with $\mu_{ggF,\gamma\gamma},\mu_{ggF,ZZ}$ focus
@@ VI.4 Convergence Metrics
Figure 12: Total loss as function of generation. Only the first 500 generations are shown. The solid (broken) lines are for the sampler CMAES (Random) while the color green (blue) indicates whether HiggsTools was used (not used) in the loop. The random sampler curves are over random generations of $12$ poitns, the same population size as CMAES. The shaded regions represent $0.95$ confidence intervals computed using a bootstrap of 100 runs.
@@
Despite the suggestive previous plot, not all scans converge within the budget, which we set to 100 runs of 2000 generations for each case in table 1 without HiggsTools, and to 200 runs of 2000 generations for the cases with HiggsTools. [^12] One could alternatively increase the budget for the HiggsTools by increasing the number of generations to 4000. Intuitively, more runs provide a more global picture, whereas longer runs allow for longer explorations of the valid region. The choice between allocating more budget to one over the other depends on the intended study. We present these metrics in table 2, where in the last two columns we show the fraction of the obtained valid points that are within each of the alignment cases, AL-1 from eq. 27 and AL-2 from eq. 28. We see that, while most points are within AL-1, only a minority are within AL-2. More interestingly, we observe how the scan with novelty reward in the $\alpha_{1},\ \beta_{1},\ \alpha_{2},\ \beta_{2}$ subspace of the parameter space has produced the most points away from the alignment limit. This can be visually understood in fig. 1 and fig. 2 where it is clear that the novelty reward is guiding CMAES to values of $\alpha_{1},\ \beta_{1}$ that are beyond the alignment limit bounds. This is a feature of the versatility of our methodology, as we can perform dedicated scans explicitly away from previously considered priors and regions of the parameter space.
@@
Sampling Scan Converged Runs Within AL-1 Within AL-2 CMAES No penalty 97 out of 100 $1.00$ $1.5\times 10^{-2}$ Parameter novelty reward 95 out of 100 $0.90$ $5.1\times 10^{-3}$ Observable novelty reward 90 out of 100 $0.99$ $1.8\times 10^{-4}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus 94 out of 100 $0.85$ $5.9\times 10^{-3}$ $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ focus 91 out of 100 $1.00$ $4.2\times 10^{-5}$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ focus 90 out of 100 $0.95$ $6.6\times 10^{-3}$ CMAES Parameter novelty reward 92 out of 200 $0.92$ $1.1\times 10^{-2}$ w/ HiggsTools Observable novelty reward 111 out of 200 $0.94$ $1.4\times 10^{-2}$ $\alpha_{1}$, $\beta_{1}$, $\alpha_{2}$, $\beta_{2}$ focus 102 out of 200 $0.84$ $4.1\times 10^{-2}$ $\mu_{ggF,\gamma\gamma}$, $\mu_{ggF,Z\gamma}$ focus 101 out of 200 $0.93$ $4.1\times 10^{-3}$ $m_{H_{1}^{+}}$, $m_{H_{2}^{+}}$ focus 91 out of 200 $0.98$ $5.3\times 10^{-2}$
@@
Table 2: Convergence and coverage statistics of the different scans presented in table 1. The last two columns represent the fractions of valid points obtained by each of the scans that are within both alignment limits. While the first CMAES rows correspond to scans without HiggsTools in the loop, the corresponding fractions of points in both alignment cases are computed using points that have passed HiggsBounds constraints after the scan.
@@
Figure 13: Number of valid points founds as function of generation. Only the first 500 generations are shown. The solid (broken) lines are for the sampler CMAES (Random) while the color green (blue) indicates whether HiggsTools was used (not used) in the loop. The random sampler curves are over random generations of $12$ points, the same population size as CMAES. The shaded regions represent $0.95$ confidence intervals computed using a bootstrap of 100 runs.
Before HiggsTools After HiggsTools Points Efficiency Points Efficiency AL-1 21 in $4.4\times 10^{12}$ $\mathcal{O}(10^{-11})$ 0 in $4.4\times 10^{12}$ $<\mathcal{O}(10^{-12})$ AL-2 13 701 in $10^{10}$ $\mathcal{O}(10^{-6})$ 510 in ×$\times 10^{10}$ $\mathcal{O}(10^{-8})$ Random 23 in $10^{13}$ $\mathcal{O}(10^{-12})$ 1 in $10^{13}$ $\mathcal{O}(10^{-13})$
@@
In fig. 14 we use a boxplot to represent the distribution of the value of the generation number when the first valid point was obtained for the runs with and without HiggsTools in the loop. We see that checking for HiggsBounds postpones the discovery of the first good point by a factor of around $2$ in terms of generation number. However, this is far better than with random sampling (both purely random and around the alignment limit) where only at most $5\%$ of the points that respect all other constraints survive HiggsBounds constraints, table 3. Additionally, we note that the number of trial points needed to find a good valid point is around $\mathcal{O}(250-750)\times 12$ when not including HiggsBounds constraints in the loop, and $\mathcal{O}(400-1300)\times 12$ when including HiggsBounds constraints in the loop. Recall that the efficient sampling of the random sampler is (estimated) to be $\mathcal{O}(10^{-12})$ and $\mathcal{O}(10^{-13})$ respectively, which means that our methodology improves the sampling efficiency by a factor of 8 orders of magnitude, even if we do not consider the near perfect sampling efficiency after convergence during the density penalty-guided exploration phase. Unsurprisingly, the improvement over the AL-2 alignment limit sampling strategy is more modest, with around four orders of magnitude when considering HiggsBounds constraints, but we reiterate that this only pertains to the first convergence and that our algorithm can then continue to explore the region found and go beyond the alignment limit bounds, as was highlighted in the previous section.